A person of mass 77 kg stands at the center of a rotating merry-go-round platform of radius 2.8 mand moment of inertia 840 kg⋅m2 . The platform rotates without friction with angular velocity 0.95 rad/s . The person walks radially to the edge of the platform. Calculate the rotational kinetic energy of the system of platform plus person before and after the person's walk.
Answer #1

Here,

mass , m1 77 Kg

r = 2.8 m

I = 840 Kg.m^2

wi = 0.95 rad/s

before the walk

rotational energy of system = 0.5 * I * wi^2

rotational energy of system = 0.5 * 840 * 0.95^2

rotational energy of system = 379.1 J

-----

let the final angular speed is wf

Using consevration of angular momentum

( 77 * 2.8^2 + 840 ) * wf = 840 * 0.95

wf = 0.604 rad/s

after the walk

rotational energy of system = 0.5 * I * wi^2

rotational energy of system = 0.5 *( 77 * 2.8^2 + 840 ) * 0.604^2

rotational energy of system = 241 J

the rotational energy of system after the walk is 241 J